The market for gravel has been estimated to have the following supply and demand relationships:

P = 100 - .01 * QD
P = 10 + .01 * QS
Calculate the equilibrium price and quantity.
Answer #1

ANSWER:

The equilibrium price can be found by equating S to D in terms of Q.

10 + 0.01q = 100 - 0.01q

0.01q + 0.01q = 100 - 10

0.02q = 90

q = 90 / 0.02

q = 4500

P = 10 + 0.01q

P = 10 + 0.01 * ( 4500)

P = 10 + 45

P = 55

The price is $55 per ton.